Grade 12Algebra

A rectangular sheet of paper is 12 ½ cm long and 10 whole 2/3 cm wide. Find its perimeter.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

From the question, it is given that,
Length = 12 whole 1⁄2 cm
= 25/2 cm
Breadth = 10 whole 2/3 cm
= 32/3 cm
We know that,
Perimeter of the rectangle = 2 × (length + breadth)
= 2 × [(25/2) + (32/3)]
= 2 × {[(25 × 3) + (32 × 2)]/6}
= 2 × [(75 + 64)/6]
= 2 × [139/6]
= 139/3 cm
Hence, the perimeter of the sheet of paper is 46 whole 1/3 cm

Thanks